Jack is 3 times as old as lacy. 5 years ago the sum of their ages was 30. How old are they now?

Answer by macston(5194)

.
L=Lacy's age; J=Jack's age=3L
(L-5)+(J-5)=30
(L-5)+(3L-5)=30
4L-10=30
4L=40
L=10
ANSWER 1: Lacy is 10 years old.
J=3L=3(10yrs)=30 yrs
ANSWER 2: Jack is 30 years old.